# Rethinking Cross-Attention for Infrared and Visible Image Fusion

### v1

@@ IV-F Computational Efficiency
The average computation time of seven different fusion methods across four IV datasets is listed in Table VII. The most efficient and second most efficient methods in each dataset are highlighted in bold and underlined, respectively. It can be seen that U2Fusion performs the highest efficiency while our method and DenseFuse both rank second. This is because our ATFuse method consumes time to alternately extract common information. Among all fusion methods, it is apparent that our ATFuse is powerful and relatively lightweight in terms of fusion performance and running speed over others.
V CONCLUSION
@@
In this article, we observed that using the cross-attention mechanism directly for multimodal image fusion only extracted common information between source images, while ignoring the extraction of discrepancy information. Therefore, we modified cross-attention and proposed DIIM to extract discrepancy information from two modality images. Additionally, to mine and integrate long dependencies, we designed an alternate common information injection module (ACIIM) based on the vanilla cross-attention structure. Based on these two modules, we developed a novel IV-IF network named ATFuse. Specifically, we first applied the DIIM module to extract discrepancy information between source images, and then alternately extracted common information. To realize a good balance between salient information and detail information retention, we proposed a segmented pixel loss function, which utilized different constraint conditions for diverse pixel parts according to the importance of pixels. The proposed method has been validated for its effectiveness on the RoadScence and MSRS datasets. Furthermore, we applied this method to other IV datasets, indicating that our method has excellent generalization ability. In comparison with six state-of-the-art fusion methods, extensive experiments show that ATFuse achieves the best fusion efficiency in both evaluations and computation burden.

### v2

@@ 4.6 Computational Efficiency
The average computation time of nine different fusion methods across all IV datasets is listed in Tab. 7. The most efficient and second most efficient methods are highlighted in bold and underlined, respectively. It can be seen that U2Fusion performs the highest efficiency while DenseFuse ranks second. The proposed ATFusion ranks third. This is because our method consumes time to alternately extract common information. Among all fusion methods, it is clear that our ATFusion is powerful and relatively lightweight in terms of fusion performance and running speed over others.
Table 7: Average running time of one image pair by using different fusion methods (unit: seconds). The best and the second best performance is marked by bold and underlined, respectively.
Methods DenseFuse U2Fusion RFN-Nest SwinFuse SwinFusion AEFusion DATFuse LRRNet Ours
Time (s) 0.0067 0.0056 0.2562 0.6888 2.0278 0.2945 0.0463 0.6052 0.0080
Figure 14: A failure case. Brightness degradation appears in fusion results.
5 Conclusion
@@
In this paper, we observed that using the cross-attention mechanism for multimodal image fusion only extracted common information between the two source images, while ignoring the extraction of discrepancy information. Therefore, we modified cross-attention and proposed DIIM to extract discrepancy information from two modality images. Additionally, to mine and integrate long dependencies, we designed an alternate common information injection module (ACIIM) based on the vanilla cross-attention structure. Based on these two modules, we developed a novel VIF network named ATFusion. Specifically, we first applied the DIIM module to extract discrepancy information between source images, and then alternately extracted common information. To realize a good balance between salient information and detail information retention, we proposed a segmented pixel loss function, which utilized different constraint conditions for diverse pixel parts according to the importance of pixels. The proposed method has been validated for its effectiveness on the MSRS dataset. Furthermore, we applied this method to the RoadScene and RGB-NIR Scene datasets, indicating that our method has excellent generalization ability. In comparison with eight state-of-the-art fusion methods, extensive experiments show that ATFusion achieves the best fusion efficiency in both quality assessment and computation burden.
A limitation of our method is the poor ability to handle low-light scene information of the source images. Fig. 14 shows a representative case to illustrate this phenomenon. As can be seen, our method struggles to improve the quality degradation problem in night scenes. A possible reason for this visual performance is that our method tends to fuse the discrepancy information of the two images. However, when the source images all face dark conditions and are difficult to distinguish, our approach has difficulty in effectively extracting their discrepancy information. In future work, we will aim to design a low-light enhancement module to address the brightness degradation problem in VIF tasks.
[…]